OpenStudy (anonymous):

is log (8 * 5) = (log8) * (log5) true or false?

OpenStudy (anonymous):

false

OpenStudy (anonymous):

should be \[\log(8\times 5)=\log(8)+\log(5)\]

OpenStudy (bahrom7893):

yup Log(A*B) = Log(A)+Log(B)